If a 300.0 mL sample of a gas is heated at constant pressure from 25.0ºC to 55.0ºC, its new volume is
_____
_____
mL?
1
Expert's answer
2016-10-25T14:56:11-0400
Write the ideal gas law:
PV=nRT, where
P is the pressure of the gas,
V is the volume of the gas,
n is the number of moles of gas,
R is the universal gas constant,
T is the absolute temperature of the gas.

In our case pressure and number of moles is constant. So transform the equation and write it
for the initial (indexed with “0”) and final (indexed with “1”) state of the system:

V/T=nR/P = constant.

V1/T1 = V0/T0.

Therefore V1 = V0*T1/T0.

Note that T is measured in kelvins. To convert from oC to kelvins we have to add 273.2.
Thus T0 = 25.0+273.2=298.2 K; T1 = 55.0+273.2 = 328.2 K.

Calculations:
V1 = 300.0mL*328.2K/273.2K = 360.4 mL